OpenStudy (anonymous):

Solve the triangle. A = 52°, b = 10, c = 7

OpenStudy (anonymous):

@Hero @zepdrix

zepdrix (zepdrix):

Drawing a picture might help

OpenStudy (anonymous):

ur right I kept drawing it as a right triangle so that made a huge difference

zepdrix (zepdrix):

oh hehe

zepdrix (zepdrix):

Looks like we have to apply the Law of Cosines here. I don't think we have enough information for Law of Sines.

OpenStudy (anonymous):

what is the law again?

zepdrix (zepdrix):

\[\Large\rm a^2=b^2+c^2-2bc cosA\]I've written it a little different here, because we want to use angle A.

zepdrix (zepdrix):

Normally it looks like:\[\Large\rm c^2=a^2+b^2-2ab~\cos C\]

OpenStudy (anonymous):

oh yeah I remember! So does that make a=12?

zepdrix (zepdrix):

i dunno, im not that fast lol.. hold on XD

OpenStudy (anonymous):

LOL sorry

zepdrix (zepdrix):

Hmm, no I'm not coming up with 12. Let's see what's going on.

OpenStudy (anonymous):

i did it one time and i got 7.9 but i thought that was wrong could that be it? or am i completely off?

zepdrix (zepdrix):

\[\Large\rm a^2=b^2+c^2-2bc~cosA\]\[\Large\rm a^2=10^2+7^2-2(10)(7)~\cos(52^o)\]

zepdrix (zepdrix):

\[\Large\rm a^2=149-140\cos(52^o)\]\[\Large\rm a^2=149-86.2\]\[\Large\rm a^2=62.81\]

zepdrix (zepdrix):

Ya it looks like the 7 point something is what we were looking for.

zepdrix (zepdrix):

Taking the square root, I'm coming up with something like 7.925

OpenStudy (anonymous):

yay ok is C=43.9?

OpenStudy (anonymous):

it keeps feeling wrong

zepdrix (zepdrix):

Law of Sines gives us:\[\Large\rm \frac{\sin(52^o)}{7.93}=\frac{\sin(C)}{7}\]Solving that gives ussssss, Yah looks like about 44 :)

OpenStudy (anonymous):

cool thank you so much I can find the next one myself! :)

zepdrix (zepdrix):

cool c: